# µGUIDE: a framework for microstructure imaging via generalized uncertainty-driven inference using deep learning

### v1

@@ 3.4 Biophysical Models
The extended SANDI model requires for the signal fractions to sum to 1, that is $f_{n}+f_{s}+f_{e}=1$. We define new parameters $k_{1}$ and $k_{2}$, uniformly sampled between 0 and 1, and use the following equations to get the corresponding signal fractions:
@@
To ensure comparability of results, we extract the same number of features $N_{f}$ using the MLP as the number of summary statistics for each model. We therefore use $N_{f}=[7,6,6]$ for the Ball&Stick, the SM and the extended-SANDI model, respectively.
3.5 Validation in Numerical Simulations
We start by validating the proposed method by comparing the obtained posterior distributions with a classical method for posterior estimations (MCMC) on the Standard Model. We generated simulations following the same acquisition protocol as the real data (see section 3.6), and added a Rician noise with a Signal-to-Noise Ratio (SNR) of 50. We then estimated the posterior distributions using both µGUIDE and the MDT toolbox (Harms and Roebroeck, ) to perform the MCMC computations.
We then compared the proposed framework to a state-of-the-art method for posterior estimation (Jallais et al., ). This method relies on manually-defined summary statistics, while µGUIDE automatically extracts some features using an embedded neural network. µGUIDE was trained directly on noisy simulations. The manually-defined summary statistics were extracted from these simulations and then used as training dataset for a MAF, similarly to (Jallais et al., ).
At least, we show that µGUIDE can be applicable to any model. While the complexity of the models increase, more degeneracies can be found. The degeneracies are inherent to the model definition, and are not induced by the noise. µGUIDE allows to highlight those degeneracies and quantify the confidence in the obtained estimates.
The training was performed on $N=10^{6}$ numerical simulations for each model, computed using the MISST package (Ianuş et al., ) and random combinations of the model parameters, each uniformly sampled from the previously defined ranges. An open-source package for µGUIDE is available at: https://github.com/mjallais/uGUIDE. It relies on the pyro (Bingham et al., ) package for carrying the probabilistic programming. Both CPU and GPU are supported.
3.6 dMRI Data Acquisition and Processing
@@ 3.5 Validation in Numerical Simulations
Data were preprocessed using tools from the FSL (Andersson et al., ; Andersson and Sotiropoulos, )) and MRTrix (Veraart et al., b; Vos et al., ) software packages. The preprossing steps included denoising (Veraart et al., b), drift correction (Vos et al., ), eddy correction (Andersson and Sotiropoulos, ), correction for gradient non-linearity distortions (Glasser et al., ) with spatio-temporal b-value/vector tracking (Rudrapatna et al., ), and Gibbs ringing artefacts correction (Kellner et al., ).
3.7 dMRI Data Analysis
Diffusion signals were first normalized by the mean non-diffusion-weighted signals acquired for each voxel. Each voxel was then estimated in parallel using the µGUIDE framework. For each observed signal $\boldsymbol{x_{v}}$ (i.e. for each voxel), we drew 50000 samples via rejection-sampling from $q_{\phi}(\theta_{i}|\boldsymbol{x_{v}})$ for each model parameter $\theta_{i}$, allowing to retrieve the full posterior distributions. We fitted two Gaussian distributions to the obtained posterior distributions to check whether a voxel was considered as degenerate. A voxel is considered as degenerate if the derivative of the fitted Gaussian distributions changes signs more than once (i.e. multiple local maxima), and if the two Gaussian distributions are not overlapping (the distance between the two Gaussian means is inferior to the sum of their standard deviations). If a posterior distribution was not deemed degenerate, the maximum-a-posteriori, uncertainty and ambiguity measures were extracted from the posterior distributions.
@@
Computations have been performed both on CPU and GPU (NVIDIA GeForce RTX 4090).
4 Results
4.1 Simulations
4.1.1 Comparison with MCMC
We performed a comparison between the posterior distributions obtained using µGUIDE and MCMC, a classical Bayesian method. Figure 3A shows posterior distributions on two exemplar simulations with $\text{SNR}=50$ using the SM, obtained with 15000 samples. Sharper and less biased posterior estimations are obtained using µGUIDE. Figure 3B presents histograms for each model parameter of the bias between the ground truth value used to simulate a signal, and the maximum-a-posteriori of the posterior distributions obtained with either µGUIDE or MCMC, on 200 simulations. Results indicate that the bias is similar or smaller using µGUIDE. Overall, µGUIDE posterior distributions are more accurate than the ones obtained with MCMC.
[…]
Figure 3: Comparison between µGUIDE and MCMC. A) Posterior distributions obtained using either µGUIDE or MCMC on two exemplar simulations with the SM ($\text{SNR}=50$). B) Bias between the ground truth values used for simulating the diffusion signals, and the maximum-a-posteriori extracted from the posterior distributions using either µGUIDE or MCMC. Sharper and less biased posterior distributions are obtained using µGUIDE.
4.1.2 The importance of feature selection
Figure 4 shows the maximum-a-posteriori extracted from the posterior distributions versus the ground truth parameters used to generate the diffusion signal with the Ball&Stick, SM and extended-SANDI models. Results are shown on 100 noise-free simulations with random parameter combinations. The orange dots have been obtained using µGUIDE, that is using a MLP for data reduction and extracting respectively $N_{f}=[7,6,6]$ features in each model. These features are then passed to the normalizing flow for the posterior density estimation. The blue dots are obtained using manually-defined summary statistics, i.e. the direction-averaged signal for the Ball&Stick model, the LEMONADE system of equations for the SM (Novikov et al., b) and the summary statistics defined in (Jallais et al., ) for the extended-SANDI model. Less biased MAPs with lower ambiguities and uncertainties are obtained with µGUIDE, indicating that the MLP allows for the extraction of additional information not contained in the summary statistics, helping to solve the inverse problem. µGUIDE generalizes the method developed in (Jallais et al., ) to make it applicable it to any forward model and any acquisition protocol, while making the estimates more precise and accurate thanks to the automatic feature extraction.
Figure 4: Maximum-A-Posterioris (MAP) extracted from the posterior distributions vs ground truth parameters used for generating the signal for the Ball&Stick, SM and extended SANDI models. Orange points correspond to the MAPs obtained using MLP-extracted features (µGUIDE) and the blue ones to the MAPs with the manually-defined summary statistics. The summary statistics used in those three models are the direction-averaged signal for the Ball&Stick model, the LEMONADE system of equations (Novikov et al., b) for the SM, and the summary statistics defined in (Jallais et al., ) for the extended-SANDI model. Results are shown on 100 exemplar noise-free simulations with random parameter combinations. The optimal features extracted by the MLP allow to reduce the bias and variance of the obtained microstructure posterior distributions.
4.1.3 µGUIDE highlights degeneracies
Figure 5 presents the posterior distributions of microstructure parameters for the Ball&Stick, Standard Model and the extended-SANDI models obtained with µGUIDE on exemplar noise-free simulations. Blue curves correspond to non-degenerate posterior distributions, while the red ones present at least one degeneracy for one of the parameters. As the complexity of the model increases, degeneracies in the model definitions appear. This figure showcases µGUIDE ability to highlight those degeneracies in the model parameter estimation. The degeneracies are inherent to the model definition, and are not an artefact generated by noise.
Figure 5: Posterior distributions (diagonal) and joint posterior distributions (upper-diagonal) of the microstructure parameters for the Ball&Stick, SM and extended SANDI models, obtained using µGUIDE on exemplar noise-free simulations. As the complexity of the model increases, degeneracies appear. µGUIDE allows to highlight those degeneracies present in the model definition.
Tables 1 and 2 present the number of degeneracies for each parameter in the three models, on 10000 simulations. Table 1 considers noise-free simulations and the training and estimations were performed on CPU. Table 2 reports results on noisy simulations (Rician noise with $\text{SNR}=50$), with training and testing performed on a GPU. The time needed for the inference and to estimate the posterior distributions on 10000 simulations, define if they are degenerate or not, and extract the MAP, uncertainty, ambiguity, are also reported.
Table 1: Number of degeneracies per parameter on 10000 noise-free simulations. Training and estimations of the posterior distributions was performed on CPU. Time for training each model and time for estimating posterior distributions of 10000 noise-free simulations, define if they are degenerate or not, and extract the MAP, uncertainty, ambiguity are also reported.
Table 2: Number of degeneracies per parameter on 10000 noisy simulations(Rician noise with $\text{SNR}=50$). Training and estimations of the posterior distributions was performed using a GPU. Time for training each model and time for estimating posterior distributions of 10000 noisy simulations, define if they are degenerate or not, and extract the MAP, uncertainty, ambiguity are also reported.
4.2 Application to real data
After demonstrating that the proposed framework provides good estimates in the case of simulations, we applied µGUIDE to both a healthy volunteer and a participant with epilepsy. The estimation of the posterior distributions is done independently for each voxel. To easily assess the values and the quality of the fitting, we are plotting the maximum-a-posterior, ambiguity and uncertainty maps, but the posterior distributions are available for all the voxels. Voxels presenting a degeneracy are highlighted with a red dot.
4.2.1 Healthy volunteer
We applied µGUIDE to a healthy volunteer, using the Ball&Stick, SM and extended-SANDI models. Figure 6 presents the parametric maps of an exemplar set of model parameters for each model, alongside their degeneracy, uncertainty and ambiguity. The Ball&Stick model presents no degeneracies, the SM presents some degeneracies, and the extended-SANDI model a lot more. This result is expected, as the complexity of the models increases, leading to more combinations of tissue parameters that can explain an observed signal. Measures of ambiguity and uncertainty allow to quantify the confidence in the estimates and help interpreting the results.
Figure 6: Parametric maps of the Ball&Stick (top), SM (middle) and extended-SANDI model (bottom), obtained using the µGUIDE framework: maximum-a-posteriori estimation (MAP), uncertainty and ambiguity measures, overlayed with voxels considered degenerate (red dots).
4.2.2 Participant with epilepsy
Figure 7 demonstrates µGUIDE application to a participant with epilepsy, using the SM. Noteworthy, the axonal signal fraction estimates within the epileptic lesion show low uncertainty and ambiguity measures hence high confidence, while orientation dispersion index estimates show high uncertainty and ambiguity suggesting low confidence, cautioning the interpretation. The two model parameters also show degeneracy in different regions of the lesion.
Figure 7: Parametric maps of a participant with epilepsy obtained using the µGUIDE framework with the SM, superimposed with the grey matter (black) and white matter (white) lesions segmentation. First quartile, mean value and third quartile of the maximum-a-posterior, uncertainty and ambiguity measures are reported in the two regions of interest. Lower MAP values are obtained in the lesions for the axonal signal fraction and the orientation dispersion index compared to healthy tissue. Higher uncertainty and ambiguity ODI values are reported, suggesting less stable estimations.
5 Discussion
5.1 Applicability of µGUIDE to multiple models
The µGUIDE framework offers the advantage of being easily applicable to various biophysical models and representations, thanks to its data-driven approach for data reduction. The need to manually define specific summary statistics that capture the relevant information for microstructure estimation from the multi-shell diffusion signal is removed. This also eliminates the acquisition constraints that were previously imposed by the summary statistics definition. The extracted features contain additional information compared to the summary statistics, resulting in a notable reduction in bias and variance in the estimated posterior distributions. Consequently, µGUIDE improves parameters estimation over current methods (e.g. Jallais et al. ).
In this study, we presented applications of µGUIDE to brain microstructure estimation using three well-established biophysical models, with increased complexity: the Ball&Stick model, the Standard Model, and an extended-SANDI model. However, our approach is not limited to brain tissue and can be extended to different organs by employing their respective favored models, such as VERDICT in prostate imaging (Panagiotaki et al., b), or even adapted to different imaging modalities. This versatility underscores the broad applicability of our proposed approach across various biological systems and imaging techniques.
It is important to note that µGUIDE is still a model-dependent method, meaning that the training process is based on the specific model being used. Additionally, the number of features extracted by the MLP needs to be predetermined. One way to determine the number of features is by matching it with the number of parameters being estimated. Alternatively, a dimensionality-reduction study using techniques like tSNE can be conducted to determine the optimal number of features.
5.2 Quantifying confidence to guide interpretation
Quantifying confidence in an estimate is of crucial importance. As demonstrated by our pathological example, changes in the tissue microstructure parameters can help clinicians decide which parameters are the most reliable and better interpret microstructure changes within diseased tissue.
Variance observed in the posterior distributions can be attributed to several factors. The presence of noise in the signal contributes to irreducible variance, decreasing the confidence in the estimates as the noise level increases. Another source of variance can arise from the choice of acquisition parameters. Different acquisitions may provide varying levels of confidence in the parameter estimates. Under-sampled acquisitions or inadequate b-shells may fail to capture essential information about a tissue microstructure, such as soma or neurite radii, resulting in inaccurate estimates. Simulations can guide users in determining whether an acquisition is suitable for estimating parameters of a given model.
Point estimate methods only provide one set of tissue parameters that could explain an observed signal, hiding model degeneracies. To mitigate the issue of degeneracy, several studies have focused on constraining the inverse problem in order to eliminate or minimize degeneracies (e.g. Fieremans et al. (a); Zhang et al. ). One popular approach is NODDI (Neurite Orientation Dispersion and Density Imaging) (Zhang et al., ), which attempts to stabilize the solution by imposing constraints on model parameters. However, those constraints may be not biologically plausible in some conditions (Novikov et al., b). As a consequence, a bias may be introduced in the parameter estimation and the inverse problem may remain largely degenerate (Jelescu et al., c). We propose instead to highlight the degeneracy via the estimation of the posterior distributions. The user can then make an informed choice and decide on the plausibility of the proposed solutions.
The presence of degeneracy in the solution of the inverse problem is influenced by the complexity of the model being used. In recent years, researchers have introduced increasingly sophisticated models to better represent brain tissue, such as NEXI (Jelescu et al., a) and eSANDIX (Olesen et al., ), that take into account exchanges between neurites and the extra-cellular space. By applying µGUIDE, it becomes possible to gain insights into the degree of degeneracy within a model and to assess the balance between model realism and the ability to accurately invert the problem.
5.3 An efficient framework
One notable advantage of the proposed framework is its amortized nature. With this approach, the training process is performed only once, and thereafter, the posterior estimations can be independently obtained for all voxels. This amortization enables efficient estimations of the posterior distributions. In fact, µGUIDE outperforms in terms of speed conventional Bayesian inference methods such as MCMC. The time savings achieved with µGUIDE make it a highly efficient and practical tool for estimating posterior distributions in a timely manner.
6 Conclusion
We proposed a general Bayesian framework, dubbed µGUIDE, to efficiently estimate posterior distributions of tissue microstructure parameters. For any given acquisition and signal model/representation, µGUIDE improves parameters estimation over existing methods. It allows to highlight existing degeneracies, and quantify confidence in the estimates, guiding results interpretation.
Acknowledgment
This work, MJ and MP are supported by UKRI Future Leaders Fellowship (MR/T020296/2). We are thankful to Dr. Dmitri Sastin and Dr. Khalid Hamandi for sharing their dataset from a participant with epilepsy.

### v3

@@ 4.3 Application of µGUIDE to biophysical modelling of dMRI data
The extended-SANDI model requires for the signal fractions to sum to 1, that is $f_{n}+f_{s}+f_{e}=1$. To uniformly cover the simplex $f_{n}+f_{s}+f_{e}=1$, we define two new parameters $k_{1}$ and $k_{2}$, uniformly sampled between 0 and 1, and use the following equations to get the corresponding signal fractions:
@@
To ensure comparability of results, we extract the same number of features $N_{f}$ using the MLP as the number of summary statistics for each model. We therefore use $N_{f}=6$ for the Ball&Stick, the SM and the extended-SANDI models.
4.3.2 Validation in Numerical Simulations
We start by validating the proposed method using posterior predictive checks (PPC) and simulated signals from model 2 (see more details in Appendix C). Since PPC alone does not guarantee the correctness of the estimated posteriors, we further validated the obtained posterior distributions comparing them with the Adaptive Metropolis-Within-Gibbs MCMC (Roberts and Rosenthal, 2009). We generated simulations following the same acquisition protocol as the real data (see Section 4.3.3), and added Rician distributed noise with a Signal-to-Noise-Ratio (SNR) of 50. We then estimated the posterior distributions using both µGUIDE and the MCMC method implemented in the MDT toolbox (Harms and Roebroeck, 2018). We initialized the sampling using a maximum likelihood estimator. We sampled 15200 samples from the distribution, the first 200 ones being used as burn-in, and no thinning. Similarly, we sampled 15000 samples from the estimated posterior distributions using µGUIDE.
Then, we show that µGUIDE can be applicable to any model. We use model 1 and 3 as examples of simpler (and non-degenerate) and more complex (and degenerate) models than model 2, respectively.
We compared the proposed framework to a state-of-the-art method for posterior estimation (Jallais et al., 2022). This method relies on manually-defined summary statistics, while µGUIDE automatically extracts some features using an embedded neural network. µGUIDE was trained directly on noisy simulations. The manually-defined summary statistics were extracted from these simulated noisy signals and then used as training dataset for a MAF, similarly to (Jallais et al., 2022).
Finally, we used µGUIDE to highlight degeneracy in all the models. While the complexity of the models increases, more degeneracy can be found. The degeneracy is inherent to the model definition, and is not induced by the noise. µGUIDE allows to highlight those degeneracies and quantify the confidence in the obtained estimates.
The training was performed on $N=10^{6}$ numerical simulations for each model, computed using the MISST package (Ianuş et al., 2017) and random combinations of the model parameters, each uniformly sampled from the previously defined ranges, with the addition of Rician distributed noise with SNR equivalent to the experimental data, i.e. 50.
4.3.3 dMRI Data Acquisition and Processing
@@ 4.3.2 Validation in Numerical Simulations
Data were preprocessed using a combination of in-house pipelines and tools from the FSL (Andersson et al., 2003; Andersson and Sotiropoulos, 2016; Smith, 2002; Smith et al., 2004) and MRTrix3 (Tournier et al., 2019) software packages. The preprocessing steps included brain extraction (Smith, 2002), denoising (Cordero-Grande et al., 2019; Veraart et al., 2016), drift correction (Vos et al., 2017; Sairanen et al., 2018), susceptibility-induced distortions (Andersson et al., 2003; Smith et al., 2004), motion and eddy current correction (Andersson and Sotiropoulos, 2016), correction for gradient non-linearity distortions (Glasser et al., 2013), and Gibbs ringing artefacts correction (Kellner et al., 2016).
4.3.4 dMRI Data Analysis
Diffusion signals were first normalized by the mean non-diffusion-weighted signals acquired for each voxel. Each voxel was then estimated in parallel using the µGUIDE framework. For each observed signal $\boldsymbol{x_{v}}$ (i.e. for each voxel), we drew 50000 samples via rejection-sampling from $q_{\phi}(\theta_{i}|\boldsymbol{x_{v}})$ for each model parameter $\theta_{i}$, allowing to retrieve the full posterior distributions. If a posterior distribution was not deemed degenerate, the maximum-a-posteriori, uncertainty and ambiguity measures were extracted from the posterior distributions.
@@
The training of µGUIDE was performed as described in 4.3.2 and an example of training dataset and input signal vector is provided in Fig. 8.
Figure 8: Example training set and input signals for µGUIDE. A) Examples of input synthetic data vectors and corresponding ground truth model parameters used in the training set of Model 1 (Ball&Stick). B) Example of input measured signals from a voxel in a healthy participant, used for inference.
All the computations were performed both on CPU and GPU (NVIDIA GeForce RTX 4090).
5 Acknowledgments
This work, Maëliss Jallais and Marco Palombo are supported by UKRI Future Leaders Fellowship (MR/T020296/2). We are thankful to Dr. Dmitri Sastin and Dr. Khalid Hamandi for sharing their dataset from a participant with epilepsy, and to Dr. Carolyn McNabb, Dr. Eirini Messaritaki and Dr. Pedro Luque Laguna for preprocessing the data of the healthy participant from the WAND data. The WAND data were acquired at the UK National Facility for In Vivo MR Imaging of Human Tissue Microstructure funded by the EPSRC (grant EP/M029778/1) and The Wolfson Foundation, and supported by a Wellcome Trust Investigator Award (096646/Z/11/Z) and a Wellcome Trust Strategic Award (104943/Z/14/Z).
Appendix
Appendix A Correlation analysis between features extracted by µGUIDE and manually-defined summary statistics
Fig. 9 presents the correlation matrices obtained from the correlation between the MLP-extracted features from µGUIDE and the manually-defined summary statistics defined in Section 4.3, considering noisy simulations (SNR=50). For each model, at least one feature extracted by µGUIDE is not or weakly correlated with the summary statistics. Additional information, not contained in the summary statistics, is extracted by the MLP from the input signal, leading to reduced bias, uncertainty and ambiguity in the parameter estimates (see Fig. 4).
Figure 9: Correlation matrices between features extracted by the MLP in µGUIDE and manually-defined summary features for the three models.
Appendix B Masked Autoregressive Flows
The conditional probability density function approximators used in this project belong to a class of neural networks called Normalizing Flows (NF) Papamakarios et al. [2021]. NFs provide a general way of transforming complex probability distributions over continuous random variables into simple base distributions $p(\boldsymbol{z)}$ (such as normal distributions) through a chain of invertible and differentiable transformations $f_{\phi}$. By applying the change of variable formula, the target distribution $q_{\phi}(\boldsymbol{\theta}|\boldsymbol{x})$ can be written as:
$q_{\phi}(\boldsymbol{\theta}\mid\boldsymbol{x})=p\left(f_{\phi}(\boldsymbol{\theta};\boldsymbol{x})\right)\left|\operatorname{det}J_{f_{\phi}}(\boldsymbol{\theta};\boldsymbol{x})\right|,$ (9)
where $\boldsymbol{z}=f_{\phi}(\boldsymbol{\theta};\boldsymbol{x})$ is invertible and differentiable (i.e. a diffeomorphism), and $J_{f_{\phi}}(\boldsymbol{\theta};\boldsymbol{x})$ is the Jacobian of $f_{\phi}(\boldsymbol{\theta};\boldsymbol{x})$. The forward direction allows for density evaluation, that is learning the mapping between the target and the base distributions, i.e. learning the parameters $\phi$. The inverse direction allows to estimate a density estimator $q_{\phi}(\boldsymbol{\theta}\mid\boldsymbol{x_{0}})$ by sampling points $\boldsymbol{z}$ from the base distribution and applying the inverse transform $f_{\phi}^{-1}(\boldsymbol{z};\boldsymbol{x_{0}})$.
A main requirement is that the flow needs to be expressive enough to approximate any arbitrarily highly complex distribution. An interesting property of diffeomorphisms is that they are closed under composition, which means that a composition of K diffeomorphisms $f=f_{1}\circ f_{2}\circ\cdots\circ f_{K}$ is also a diffeomorphism, and the Jacobian determinant is the product of the determinant of each component. Combining multiple transformations allows to increase the expressivity of the general flow. We obtain:
$\begin{array}[]{rcl}q_{\phi}(\boldsymbol{\theta}\mid\boldsymbol{x})&=&p\left(f_{\phi}(\boldsymbol{\theta};\boldsymbol{x})\right)\log\left|\operatorname{det}\prod_{k=1}^{K}J_{f_{k}}\left(\boldsymbol{\theta}_{\boldsymbol{i}};\boldsymbol{x}_{\boldsymbol{i}}\right)\right|\\[8.99994pt] &=&p\left(f_{\phi}(\boldsymbol{\theta};\boldsymbol{x})\right)\sum_{\boldsymbol{k}=\mathbf{1}}^{\boldsymbol{K}}\log\left|\operatorname{det}J_{f_{k}}\left(\boldsymbol{\theta}_{\boldsymbol{i}};\boldsymbol{x}_{\boldsymbol{i}}\right)\right|\end{array}$ (10)
Flows need to be flexible and expressive enough to model any desired distribution but also need to be computationally efficient, that is, computing the associated Jacobian determinants need to be tractable and efficient. Among a number of proposed architectures such as mixture density networks [Bishop, 1994] or neural spline flows [Durkan et al., 2019], we focused on Masked Autoregressive Flows (MAF) [Papamakarios et al., 2017], which has shown state-of-the-art performance as well as the ability to estimate multi-modal posterior distributions [Gonçalves et al., 2020, Patron et al., 2022, Papamakarios et al., 2021].
Autoregressive flows are universal approximators and have the form $z_{i}^{\prime}=\tau(z_{i};\boldsymbol{h_{i}})$ where $\boldsymbol{h_{i}}=c_{i}(\boldsymbol{z_{<i}})$ [Papamakarios et al., 2021]. $\tau$ is termed the transformer and is a stritly monotonic function parametrized by $\boldsymbol{h_{i}}$, and $c_{i}$ the $i$-th conditioner. Each $\boldsymbol{h_{i}}$ and therefore each $z_{i}^{\prime}$ can be computed independently in parallel, helping to keep a low computation time. The conditioner constraints each output to depend only on variables with dimension indices less than $i$, which makes the Jacobian of the flow lower diagonal. Its determinant can then be obtained easily as the product of its diagonal elements.
To efficiently implement the conditioner, this method relies on the Masked Autoencoder for Distribution Estimation (MADE) [Germain et al., 2015] architecture. To create a neural network that obey the autoregressive structure of the conditioner, a fully connected feedforward neural network is multiplied to binary masks, which removes some connections by assigning them a weigh of 0. The binary masks can easily be obtained by following a few simple steps (see Fig. 10 for an illustration):
  1. 1.
    Label the input and output nodes between 1 and $D$, $D$ being the dimension of the input vector $\boldsymbol{z}$.
  2. 2.
    Randomly assign each hidden unit a number between 1 and $D-1$, which indicates the number of inputs it will be connected to.
  3. 3.
    For each hidden layer, connect the hidden units to units with inferior or equal labels.
  4. 4.
    Connect output units to units with strictly inferior labels.
For µGUIDE’s implementation, we use a combination of 5 MADEs.
As a result, the MAF architecture only needs a single forward pass through the flow and, combined with the low-cost computation of the determinant, allows for fast training and evaluation of the posterior distributions.
Figure 10: Schematic of MADE autoregressive network construction.
Appendix C Posterior Predictive Checks
Posterior Predictive Checks (PPC) are a common safety check to verify inference is not wrong. The idea is to compare input signals with generated signals from samples drawn from the posterior distributions. If the inference is correct, the generated signals should look similar to the input signal.
We performed the following steps:
  • Sample $N$ $\boldsymbol{\theta_{i}}$ from the prior distribution: $\boldsymbol{\theta_{i}}\sim p(\boldsymbol{\theta})$
  • Generate the corresponding signals using the forward model: $\boldsymbol{x_{i}}=\mathcal{M}(\boldsymbol{\theta_{i}})$
  • Perform the inference and estimate the posterior distributions $p(\boldsymbol{\theta_{i}}|\boldsymbol{x_{i}})$
  • Sample $N_{PP}$ samples $\boldsymbol{\theta_{i,s}}$ from $p(\boldsymbol{\theta_{i}}|\boldsymbol{x_{i}})$
  • Reconstruct the signals from the sampled $\boldsymbol{\theta_{i,s}}$ using the forward model: $\boldsymbol{x_{i,s}}=\mathcal{M}(\boldsymbol{\theta_{i,s}})$
  • Compare the obtained $\boldsymbol{x_{i,s}}$ with $\boldsymbol{x_{i}}$.
Fig. 11 presents results on model 2 (SM), on both noise-free and noisy signals (Rician noise with SNR=50) for $N=10$ random combinations of model parameters, and $N_{PP}=100$. As dMRI data have a high dimensionality, we report the direction-average signal. Plain lines show the signals $\boldsymbol{x_{i}}$, and the shaded areas correspond to the area in which the corresponding $\boldsymbol{x_{i,s}}$ fall. $\boldsymbol{x_{i}}$ lie within the support of $\boldsymbol{x_{i,s}}$, indicating the inference is not wrong. Note that the support of $\boldsymbol{x_{i,s}}$ is bigger for noisy simulations, reflecting the wider posterior distributions obtained from the inference.
Figure 11: Posterior Predictive Checks. Comparison between signals $\boldsymbol{x_{i}}$ generated using random parameter combinations and their reconstructions using samples from $p(\boldsymbol{\theta_{i}}|\boldsymbol{x_{i}})$.